Assume that you contribute $340 per month to a retirement plan for 25 years. Then you are able to increase the contribution to $540 per month for the next 25 years. Given a 7.2 percent interest rate, what is the value of your retirement plan after the 50 years? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Find Future value of multiple annuities

Please answer correctly. THANK YOU!
Answer #1

Calculating Value after 25 years,

Using TVM Calculation,

FV = [PV = 0, PMT = 340, I = 0.072/12, N = 300]

FV = $284,307.82

Calculating Value after 50 years,

Using TVM Calculation,

FV = [PV = 284,307.82, PMT = 540, I = 0.072/12, N = 300]

FV = $2,162,283.82